OpenStudy (anonymous):

Simplify this expression: x^2+7x+10/x+2 divided by x^2+2x-15/2x+4

OpenStudy (e.mccormick):

Do you know what a fraction over a fraction is the same as? It relates to multiplication.

OpenStudy (e.mccormick):

\[\frac{\frac{x^2+7x+10}{x+2}}{\frac{x^2+2x-15}{2x+4}}\]

OpenStudy (anonymous):

x-3/2x-4

OpenStudy (goldphenoix):

\[\frac{ x^2+7x+10/x+2 }{ x^2+2x-15/2x+4 }\]

OpenStudy (anonymous):

@e.mccormick I got x-3/2x+4

OpenStudy (e.mccormick):

I am not sure you have that right. The lower fraction gets inverted for the cancels, so how would the 2x+4 stay on the bottom?

OpenStudy (anonymous):

Oh 1/ (2x+4)(x-3)?

OpenStudy (e.mccormick):

\[\frac{\frac{x^2+7x+10}{x+2}}{\frac{x^2+2x-15}{2x+4}}\implies \frac{x^2+7x+10}{x+2}\cdot \frac{2x+4}{x^2+2x-15}\implies \\ \frac{(x+5)(x+2)}{x+2}\cdot \frac{2x+4}{(x+5)(x-3)}\implies \frac{1}{1}\cdot \frac{2x+4}{(x-3)}\]

OpenStudy (e.mccormick):

You answer was just inverted.

OpenStudy (anonymous):

Oh thank you!

OpenStudy (anonymous):

I have more problems if you dont mind helping me?

OpenStudy (e.mccormick):

Just remember that a fraction divided by a fraction is the same as the top fraction times the reciprocal of the bottom fraction. That is what lead me to the answer.

OpenStudy (anonymous):

I could really use the help with these!

OpenStudy (e.mccormick):

Do you know what the lowest common demonimatior is like for rational equations? Like the first one.... 18.

OpenStudy (anonymous):

I tried it out and I got 12x-2/(3x+1)(3x-1)

OpenStudy (e.mccormick):

How did you get the -2 on the top?

OpenStudy (anonymous):

Im sorry it was 12x-6*

OpenStudy (e.mccormick):

Do you have your work? So I can see where your making the mistake?

OpenStudy (anonymous):

Sad excuse... my left arm is fractured I cant type that much sorry

OpenStudy (anonymous):

was that right?

OpenStudy (e.mccormick):

Neither of those is right.

OpenStudy (e.mccormick):

Factor the bottom of the second fraction. Find what factor is missing in the first fraction. Multiply through by that missing factor over itself. Do NOT multiply out the bottom, but be sure to distribute the top properly. The 5 distributes to both parts of the new left fraction's top. Distribute the negative to both parts of the right fraction's top. Combined like terms. Factor. Cancel.

OpenStudy (anonymous):

I redid it agin, 4/3x-1

OpenStudy (e.mccormick):

Yes.

OpenStudy (e.mccormick):

b is correct. Let me guess, you had distributed something wrong before. =) I know because I have made that mistake myself!

OpenStudy (anonymous):

yes exactly!

OpenStudy (anonymous):

19 I got no solution

OpenStudy (e.mccormick):

So, you got 0=8 or something fun like that?

OpenStudy (e.mccormick):

Yah, it looks like a classic division by 0 error to me. So no real solutions.

OpenStudy (anonymous):

and for 18 B was 4/3x+1 not minus 1 that was A

OpenStudy (anonymous):

So for 18 was it A or B?

OpenStudy (e.mccormick):

Sorry, got called away for a min.

OpenStudy (e.mccormick):

THose notes are scribbled by now... let me rework.

OpenStudy (anonymous):

ok

OpenStudy (e.mccormick):

a. The (3x+1) cancels so (3x-1) is left.

OpenStudy (e.mccormick):

Dunno why I said b. LOL. Looking at the old notes and the new, and I canceled the same thing...

OpenStudy (anonymous):

ok so A!

OpenStudy (e.mccormick):

Yes. Absolutely.

OpenStudy (anonymous):

ok 20. I got 3 1/3 hours

OpenStudy (e.mccormick):

Not sure... I was never good at setting those up right. I know Lori is faster than Elissa. If your math supports 3 and 1/3, I would say it is good.

OpenStudy (anonymous):

what did you get for 21?

OpenStudy (anonymous):

I got D) 3

OpenStudy (e.mccormick):

Did you start with this: \[36^{(3x-1)}=6^{(4x+2)}\implies (6^2)^{(3x-1)}=6^{(4x+2)}\implies \\ 2(3x-1)=(4x+2)\]

OpenStudy (anonymous):

No, ill try again now

OpenStudy (e.mccormick):

OK, do you se what I used to get there? Exponet rules:\[\large (x^m)^n=x^{m\cdot n}\]

OpenStudy (anonymous):

Yes

OpenStudy (e.mccormick):

Then they are exponents of the same number, and they must be equal. Log rules let me drop them, set them equal, and solve.

OpenStudy (anonymous):

I got 1/2

OpenStudy (e.mccormick):

\[2(3x-1)=(4x+2)\implies 6x-2=4x+2\implies \\ 6x-2+2=4x+2+2\]Do you see your mistake that would get you 1/2 when I do it that way?

OpenStudy (e.mccormick):

You subtracted off a 2 rather than add in a 2.

OpenStudy (e.mccormick):

Or something like that.

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I got 2 now

OpenStudy (e.mccormick):

=)

OpenStudy (e.mccormick):

Now, on 22, do you know the big thing you need to be careful of?

OpenStudy (anonymous):

No what's that?

OpenStudy (e.mccormick):

\[\Large \log_b(y)=x\Leftrightarrow y=b^x\]If x is positive, can y be negative?

OpenStudy (anonymous):

No

OpenStudy (e.mccormick):

Exactly, so you need to test your answers in the originl functions.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

I tried it and got 5 then I checked and got no solution

OpenStudy (e.mccormick):

hmmm... I got something else. What did you start with?

OpenStudy (anonymous):

I meant -5

OpenStudy (e.mccormick):

I got -5 and 2. When I checked, one of those went by-by.

OpenStudy (anonymous):

2?

OpenStudy (e.mccormick):

When you put 2 into the original equation, are you ending up with a + or a - number in the log?

OpenStudy (anonymous):

I understand

OpenStudy (e.mccormick):

\[5-3x=x^2-5\implies 0=x^2-5-5+3x\implies x^2+3x-10=0\implies \\ (x-2)(x+5)=0\implies x=\{-5,2\} \\ \text{OK, so two to test.}\\ 5-3(2)\implies 5-6=-1\\ 2^2-5\implies 4-5=-1\\ 5-3(-5)\implies 5+15=20\\ (-5)^2-5\implies 25-5=20 \]Do you see what I did there? That is the entire process. You have to be careful of the small mistakes.

OpenStudy (anonymous):

23 I got A. 11

OpenStudy (e.mccormick):

Now, can I take a log, any log, of -1? 20? If yes to -1, then the answer is 2. If yes to 20, then the answer is -5. If the answer is yest to both it is both, no to both is nnone.

OpenStudy (anonymous):

I got -5,2

OpenStudy (e.mccormick):

Rememeber what I said earlier about the log of a negative number? It does not exist. Well, it is a complex number... so if you try \(\log_6(-1)\) it does not exist.

OpenStudy (e.mccormick):

That is the whole point of testing things. If you get a negative value, there is no real log.

OpenStudy (anonymous):

23 I got A. 11

OpenStudy (e.mccormick):

Yes, and I am wondering how.

OpenStudy (e.mccormick):

wait... let me make sure I did not make a mistake on one thing there.

OpenStudy (e.mccormick):

I did.. but I still did not get 11.

OpenStudy (anonymous):

Ok I'll check too

OpenStudy (e.mccormick):

WHat are you doing with the fractions in front of the logs?

OpenStudy (anonymous):

Wait I see my mistake

OpenStudy (e.mccormick):

OK.

OpenStudy (anonymous):

14

OpenStudy (e.mccormick):

yah. THat is what I got.

OpenStudy (e.mccormick):

On the last one, have they told you what base they standardly use for antilog? Because people tend to use either 10 or e, but sometimes they just say log or antilog and mean base e rather than 10!

OpenStudy (e.mccormick):

Find what base it is in this book, then 24 is a calculator problem.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

how do I find the antilogrithm?

OpenStudy (anonymous):

I got 817

OpenStudy (anonymous):

@e.mccormick is that right?